Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
OPERATION-2025
OPERATION-2024
OPERATION-2023
OPERATION-2022
OPERATION-2021
OPERATION-2020
OPERATION-2019
SASE1-2018
SASE1-2017
Experiment data
OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn
Not logged in
List
|
Message ID:
242
Entry time:
10 Jun 2023, 21:52
Author:
Jan Gruenert
Group:
PRC
Subject:
Status
The Cryo team solved the problems, beam is back.
Downtime from 17h59 until 21h03.
Attachment 1:
JG_2023-06-10_um_21.55.46.png
4.616 MB
Uploaded 10 Jun 2023, 22:58
|
Hide
|
Hide all
ELOG V3.1.4-7c3fd00